A solution contains 0.05 M acetic acid and 01 M sodium acetate. Calculate the concentration of H+,
when HC1 is added to the solution such that
[HC1] = 1.0 x 10-3M.
Kafor acetic acid = 1.8 x 10-5.
1
Expert's answer
2019-06-03T05:48:22-0400

Solution.

A solution consisting of a solution of acetic acid and sodium acetate is considered as a buffer system of a weak acid and a salt of a weak acid. Therefore, we use the following formula:


"pH = pKa - lg{\frac{C(acid)}{C(salt)}}"

pH = 5.041

If we add acid or base to the buffer, then the equilibrium will shift and the formula will take the following form:


"pH = pKa - lg{\frac{C(acid)+X}{C(salt)-X}}"

H+ + CH3COO- = CH3COOH


"pH = 4.74-lg{\frac{0.05+10^{-3}}{0.1-10^{-3}}}"

pH = 5.028


"[H+] = 9.736 \times 10^{-6}"

Answer:

"pH = 5.028"

or


"[H+] = 9.736 \times 10^{-6}"